eight boys and three girls are to sit in row for a photograph.find the probability that no two girls are together


1
Expert's answer
2020-07-13T19:53:22-0400

We are seeking the situations of the form

"(x_1\text{ boys})(\text{a girl})(x_2\text{ boys})(\text{a girl})(x_3\text{ boys})(\text{a girl})(x_4\text{ boys})",

where "x_2,x_3>0." The number of such situations is the number of sequences of nonnegative integers "x_1,x_2,x_3,x_4" such that "x_1+x_2+x_3+x_4=8" and "x_2,x_3>0" times the number of permutations between girls and boys, i.e. "8!\cdot 3!". The number of those sequences is the number of sequences of positive integers "y_1,y_2,y_3,y_4" such that "y_1+y_2+y_3+y_4=10" with the bijection "(x_1,x_2,x_3,x_4)\mapsto(x_1+1,x_2,x_3,x_4+1)=(y_1,y_2,y_3,y_4)", and we indeed have "y_1+y_2+y_3+y_4=(x_1+1)+x_2+x_3+(x_4+1)=10". It's bijective with taking a line of "10" balls and choosing "3" borders between them with the bijection

"(y_1,y_2,y_3,y_4)\mapsto (\text{borders after the balls }y_1,y_1+y_2,y_1+y_2+y_3)."

We choose "3" borders from "9" places between balls, which gives "\binom93" such sequences. Thus, the probability that no two girls sit near each other equals "\binom93\cdot 8!\cdot 3!/11!=28/55", where we divide by "11!" since it's the number of permutations of "8+3=11" boys and girls.